# PhysORD: A Neuro-Symbolic Approach for Physics-infused Motion Prediction in Off-road Driving

### v1

@@ I Introduction
Traditional physics-based models rely on motion formulas derived from fundamental principles, such as Newton’s Laws of Motion [10], for vehicle state prediction. While effective in their great generalization ability, they struggle to accurately model the complex dynamic systems and external disturbances in off-road conditions. Evolution models based on ideal assumptions, such as kinematic models [11, 12], along with state estimation algorithms like the Kalman Filter [13] with simplified noise assumptions, face difficulties in predicting long-term motion due to external disturbances.
@@ II-C Physics-infused Neural Networks
To further impose more physical property and structure constraints, such as symmetries, on the embedding space, Variational Integrator Networks (VINs) [18] utilize a structure-preserving discretization method of variational integrators to derive discrete-time motion equations that maintain both geometric structure and physical laws. Building on this, Forced Variational Integrator Networks (FVIN) extend VINs for broader system applications, incorporating energy dissipation from external controls. The LieFVIN [19] further combines VINs and FVINs to learn a discrete time symplectic approximation for robot systems evolving on the Lie group and demonstrate the learned dynamics can be used in the control of drone in simulation. However, most of their applications are limited to simple physics or simulation environments. Extending these approaches to real-world ground vehicles remains challenging.
@@ III-B Symbolic Model
We employ Lagrangian and Hamiltonian mechanics to develop a symplectic map that updates the state $s_{t}$, inspired by [18, 19]. The evolution follows the conservation law and preserves the structure imposed by the constraints on Lie manifold [28]. Hamilton mechanics describes the dynamic system in terms of energy with the generalized coordinates $q$. According to Hamilton’s Variational Principle, the actual trajectory of a system is the one where the action integral is stationary for variations over a time interval $T$:
@@ V Conclusions and Future Work
This paper introduces PhysORD, a neuro-symbolic model that infuses physics, specifically the Euler-Lagrange equation, with neural networks for motion prediction in off-road driving scenarios. By modeling the vehicle as a controlled Lagrangian system and using MLPs to estimate forces and potential energy, PhysORD addresses the challenges posed by complex dynamics and environmental uncertainties. Through extensive experiments on the TartanDrive dataset, PhysORD demonstrated superior performance over data-driven methods, achieving a significant 46.7% improvement in prediction accuracy while using 96.9% fewer parameters. This reduction in parameters notably enhances both training and inference efficiency. Moreover, PhysORD exhibits the ability to learn from limited data and generalize from short-term learning to long-term prediction. In the future, we plan to enhance the model’s performance by integrating additional environmental data, such as forward-facing images and terrain height maps, and apply it in the control and planning of real-world off-road vehicles.

### v2

@@ I Introduction
Traditional physics-based models rely on motion formulas derived from fundamental principles, such as Newton’s Laws of Motion [10], for vehicle state prediction. While effective in their generalization ability, they struggle to accurately model the complex dynamic systems and external disturbances in off-road driving. Off-road conditions introduce various challenges, such as irregular terrain with bumps and load variations, which cause the system to exhibit highly nonlinear and nonstationary behavior. Evolution models based on ideal assumptions, such as kinematic models [11, 12], along with state estimation algorithms like the Kalman Filter [13] with simplified noise assumptions, face difficulties in predicting long-term motion due to external disturbances.
@@ II-C Physics-infused Neural Networks
To further impose more physical property and structure constraints, such as symmetries, on the embedding space, Variational Integrator Networks (VINs) [18] utilize a structure-preserving discretization method of variational integrators to derive discrete-time motion equations that maintain both geometric structure and physical laws. Building on this, Forced Variational Integrator Networks (FVIN) [21] extend VINs for broader system applications, incorporating energy dissipation from external controls. The LieFVIN [19] further combines VINs and FVINs to learn a discrete time symplectic approximation for robot systems evolving on the Lie group and demonstrate the learned dynamics can be used in the control of drone in simulation. However, most of their applications are limited to simple physics or simulation environments. Extending these approaches to real-world ground vehicles remains challenging.
@@ III-B Symbolic Model
We employ Lagrangian and Hamiltonian mechanics to develop a symplectic map that updates the state $s_{t}$, inspired by [18, 19]. The evolution follows the conservation law and preserves the structure imposed by the constraints on Lie manifold [28, 29]. Hamilton mechanics describes the dynamic system in terms of energy with the generalized coordinates $q$. According to Hamilton’s Variational Principle, the actual trajectory of a system is the one where the action integral is stationary for variations over a time interval $T$:
@@ V Conclusions and Future Work
This paper introduces PhysORD, a neuro-symbolic model that infuses the Euler-Lagrange equation with neural networks for motion prediction in off-road driving scenarios. By modeling the vehicle as a controlled Lagrangian system and using MLPs to estimate forces and potential energy, PhysORD addresses the challenges posed by complex dynamics and environmental uncertainties. Experiments on the TartanDrive dataset show that PhysORD outperforms data-driven methods, improving prediction accuracy by 46.7% while using 96.9% fewer parameters. This reduction in parameters notably enhances both training and inference efficiency. Moreover, PhysORD exhibits the ability to learn from limited data and generalize from short-term learning to long-term prediction. Future work will focus on integrating environmental data, such as forward-facing images and terrain height maps, to further improve performance in real-world off-road vehicle control and planning.
Acknowledgements
This work was in part supported by the ONR award N00014-24-1-2003 and DARPA grant DARPA-PS-23-13. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of the ONR or DARPA. The authors also wish to express their gratitude for the generous gift funding provided by Cisco Systems Inc.